# A Hybrid Machine Learning Framework for Predicting Hydrogen Storage Capacities: Unsupervised Feature Learning with Deep Neural Networks

### v3

@@ Abstract
In this study, we present a sophisticated hybrid machine-learning framework that significantly improves the accuracy of predicting hydrogen storage capacities in metal hydrides. This is a critical challenge due to the scarcity of experimental data and the complexity of high-dimensional feature spaces. Our approach employs the power of unsupervised learning through the use of a state-of-the-art autoencoder. This autoencoder is trained on elemental descriptors obtained from Mendeleev software, enabling the extraction of a meaningful and lower dimensional latent space from the input data. This latent representation serves as the basis for our deep multi-layer perceptron (MLP) model, which consists of five layers and shows good precision in predicting hydrogen storage capacities. Furthermore, our results show very good agreement with the results obtained with density functional theory (DFT). In addition to addressing the limitations caused by limited and unevenly distributed data in the field of hydrogen storage materials, we also focus on discovering new materials that show promising opportunities for hydrogen storage. These materials were identified using both feature-based approaches and predictions generated by a large language model. Finally, our investigation into the effectiveness of transferring weights from the autoencoder to the MLP, in addition to the latent features, suggests that while this strategy slightly improves model performance as indicated by a slightly higher R value and lower RMSE, it emphasizes the intricate challenge of adapting pre-trained weights for specific supervised tasks.
@@
The introduction of data-driven materials design 25, 26, 27 has accelerated material discovery, processing, and manufacturing 28, 29, 30. Machine learning (ML) algorithms allow for the creation of new alloys based solely on previously collected data, either publicly available or reported in scientific literature. Metal hydrides are an ideal subject for machine learning algorithm research due to their diverse properties and broad range. Traditional methods for determining the optimized materials class and corresponding metal hydride composition based on desirable properties are difficult, time-consuming, and costly. ML learning techniques, on the other hand, enable rapid, productive, and efficient material class prediction for a specific hydrogen weight percent and operational conditions. Numerous groups have begun to use ML techniques to speed up this screening and gain more physical insight from massive amounts of data 31, 32, 33. Rezakazemi et al. used an adaptive neuro-fuzzy inference system to evaluate the performance of hydrogen-selective mixed matrix membranes under various operational conditions 34. Rahnama et al. predicted the hydrogen storage capacities in various metal hydrides using machine learning algorithms in another study and found that higher temperatures yielded higher hydrogen storage capacities 33, 35. Ahmed et al. recently predicted gravimetric and volumetric hydrogen capacities in metal-organic frameworks (MOFs) using ML algorithms 36. ML methods have been instrumental in predicting various properties of (MOFs) and even designing new structures. This approach has significantly changed the research landscape in this field, allowing for the rapid identification of promising MOF candidates 37, 38 for various applications, including hydrogen storage.
@@ 2 Methods
The architecture of our method is shown in Fig. 1, in a schematic way. We use a hybrid approach that has two parts: in the unsupervised learning part, an autoencoder is trained on the entire dataset, which consists of only unlabeled data, to learn a compact representation of the data in the bottleneck layer 39. A Multi-Layer Perceptron (MLP) is used in the supervised learning part to predict the hydrogen storage capacity. The major use of such an approach is to address the challenges of data scarcity and high feature dimensionality. In this study, the initial feature space has a high dimensionality and the available dataset is relatively small. As there are not too many machine learning studies available in the literature, it is not easy to choose proper descriptors. We have constructed 36 features using the Mendeleev software 40. To handle high-dimensional feature spaces with small datasets, we used an autoencoder to reduce the dimensionality of the data.
@@ 2.1 Unsupervised learning
Autoencoders are a kind of neural network that is used to extract features in an unsupervised manner. They are divided into two parts: an encoder, which compresses the input data into a lower-dimensional representation (also known as the latent representation or bottleneck), and a decoder, which reconstructs the original input from the compressed form 41.
@@
where is the activation function, and are the weight and bias parameters of the encoder, respectively. For the activation function, we have used the RelU function as implemented in PyTorch 42.
@@
This loss function was minimized using the Adam optimization method to find the optimal values for the parameters and say and 43. Once the autoencoder is trained, the encoder function can be used to extract features from the input data by using the activations of the lower-dimensional representation as the features.
The network architecture for both encoder and decoder includes two linear layers and dropout layers for preventing overfitting. The input to the autoencoder is first passed through a linear layer with 256 units, followed by a Leaky ReLU activation function 44. This is then followed by a dropout layer with a dropout probability of 0.3 to provide regularization.
@@ 2.2 Supervised learning
First-principles calculations were conducted employing the density functional theory (DFT) within the Vienna Ab initio Software Package (VASP) 45, 46. To determine the electronic energy, projected augmented wave (PAW) potentials and plane-wave basis sets were utilized for both core and valence electrons, employing the Perdew–Burke–Ernzerhof (PBE) functional 47, 48. The calculations accounted for long-range interactions using the semi-empirical Grimme D2 dispersion method and incorporated non-spherical contributions to the PAW potentials within the code 49. A cutoff energy of 500 eV ensured convergence of all energies. Structural optimization employed the conjugate gradient algorithm 50, with convergence criteria set at 10-7 eV for energy and 0.01 eVÅ-1 for force. Spin polarization was enabled in all calculations, except for the isolated H2 closed shell molecule.
@@ 3 Dataset
For this study, we utilized the Hydrogen Storage Materials Database shared by the HyMARC Data Hub (https://datahub.hymarc.org/). This is an openly available database that can be accessed at https://datahub.hymarc.org/en/dataset/hydrogen-storage-materials-db. This database contains information on more than 2000 hydride materials and their properties. The Hydride Database classifies the hydrides into eight classes: A2B intermetallic compounds, AB intermetallic compounds, AB2 intermetallic compounds, AB5 intermetallic compounds, complex hydrides, Mg alloys, solid solution alloys, and Misc. In the dataset, more than 950 compounds have 1-2 wt% hydrogen storage capacity (Fig. 2). The average hydrogen storage capacity of all compounds is 2.1 wt%. The dataset has as high as 20.8 wt% hydrogen storage capacity complex hydrides, which is BeB4H8 51.
@@ 5.2 Case I: Using the MLP Directly for All the Features
For our primary experiment, the MLP was trained and tested using all of the 36 features derived from the Mendeleev software. The dataset was strategically partitioned into training and testing subsets, with respective ratios of 0.8 and 0.2. The resultant outcomes of this methodology are visually represented in Fig.4. The predictive efficacy, quantified by the coefficient, was deemed unsatisfactory 52. Such results emphasize the inherent difficulties connected with managing datasets with a high number of dimensions, especially when the amount of data is small.
@@ 6 Comparison with few other supervised models
In this section, we compare the performance of two supervised models, Lasso regression and Gradient Boosting regression, with the previously mentioned hybrid approach using an Autoencoder combined with a Multilayer Perceptron (MLP) 53, 54. We evaluate the models based on their R scores, which measure the proportion of the variance in the target variable that the models can explain. The results are shown in Fig.7. The R values obtained from the Lasso regression and Gradient Boosting regression are 0.48 and 0.8, respectively. These values indicate that both models have relatively lower predictive accuracy compared to the hybrid approach. The hybrid approach achieved an R score of 0.85, outperforming both linear models. These results suggest that the hybrid approach, combining an Autoencoder with an MLP, is more effective in capturing the underlying patterns and predicting the target variable compared to the individual linear models.
@@ 6.1 Identification of new hydrogen storage materials
To find new materials for efficient hydrogen storage, we used the following approach: We calculated the Pearson correlation function for different physically meaningful features, including the target - the storage capacity (Wt), which we show in Fig. 8 55. We take the weighted sum for all nine features and compute the correlation. It can be seen that the period number and electronegativity have the highest correlation with the weight percentage. A significant Pearson correlation between the period number and electronegativity with the target hydrogen storage capacity, suggests their pivotal roles in influencing a material’s ability to store hydrogen efficiently. The period number, reflecting the electron shell structure, directly impacts its hydrogen bonding capabilities, while electronegativity, indicative of an electron attraction strength of the material, influences the stability and capacity of hydrogen storage in metal hydrides. Therefore, we first scanned the metal hydrides from the Materials Project with these two features. Next, we constructed a 36-dimensional feature space using the above materials, as before (using four statistical measures for each one). Then, we used pre-trained embeddings from the autoencoder so that the property (hydrogen weight in percentage) can be computed using the MLP. Finally, we performed DFT-based calculations to re-verify these values.
Using the above approach we identified two materials AlO and VO. Both the materials are stable from our formation energy calculations. The DFT calculated formation energies are respectively -3.34 eV/atom and -2.12 eV/atom for AlO and VO. From the DFT approach the hydrogen storage capacities were calculated using the hydrogen adsorption energies 11 shown below,
@@
To predict hydrogen storage capacity, all interstitial sites of AlO and VO are gradually filled with hydrogen, and the adoption energy is calculated 56. Figure 9 shows the change in adsorption energy with increasing hydrogen concentration. The adsorption energy increases as the hydrogen fraction increases. The maximum hydrogen fraction at which adsorption energy remains negative is used to predict hydrogen storage capacity. The predicted hydrogen storage capacities from the DFT calculations are respectively 4.61% and 3.83% while from the MLP model, we obtain values respectively 5.98% and 4.35%.
In the Fig.10, we show the density of states (DOS) for the machine learning predicted materials and their hydrides for different hydrogen loadings. It was found that AlO shows a more stable behavior when loaded with hydrogen compared to VO because VO relatively DOS at the Fermi energy for all the hydrogen loading. For both materials, H-s strongly hybridizes with the O-p states as can be seen from the projected DOS of AlOH and VOH shown in the down panel of the above figure. In the case of VOH, the anti-bonding states formed between V-d and H-s are also filled and therefore the system shows less stability.
@@ 7 Application to the Hydrogen Storage Materials Generated Using a Large Language Model
To show further the predictive capability of our approach, we apply it further to a set of materials that we generate using a large language model (LLM). Leveraging the power of GPT-2 (Generative Pre-trained Transformer version 2), a model renowned for its proficiency in natural language processing, we have generated a few new hydrogen storage materials. Unlike the resource-intensive traditional methods such as Generative Adversarial Networks (GANs) or diffusion models, which have limited testing in materials science, our approach utilizes a fine-tuned GPT-2 model, offering a more quick and targeted pathway to innovation. The details of the methodology are explained as detailed by Fu et. al 57 which leverages the advanced capabilities of LLMs originally developed for natural language processing, adapting them to the intricate language of material compositions. The fine-tuning process involves adapting a pre-trained GPT-2 model to the specific nuances of material science. This adaptation is facilitated by a custom vocabulary. which includes a comprehensive list of chemical elements and special tokens. This specialized vocabulary is crucial for accurately representing and processing material compositions. Our training process fine-tunes the model with a dataset specifically curated for hydrogen storage materials, allowing the model to learn and predict the complex patterns of material composition. The cornerstone of this methodology is the training of these models on comprehensive databases of known material compositions, including the ICSD (Inorganic Crystal Structure Database), OQMD (Open Quantum Materials Database), and Materials Project. This extensive assimilation of existing material data equips LLMs like GPT-2 with the ability to understand and replicate the intricate patterns and rules that govern material compositions. Through this process, the models are enabled to generate new, chemically plausible material compositions, which hold great potential for hydrogen storage applications. In an extension to the work by Fu et al. 57 on materials generation, we take a slightly modified approach where not only the chemical composition of materials but also a specific property: their hydrogen storage capacity is also considered to generate the chemical labels.
@@
Starting from a sequence of input tokens (comprising chemical symbols and storage capacity) the model first computes their embeddings. It then applies learned weight matrices to derive the usual queries, keys, and values matrices , , and , which are used to calculate the self-attention scores given by where is the dimensionality of the key vectors. These scores determine how much each token in the sequence should attend to every other token, capturing both local and global dependencies within the sequence. During the training phase, the objective is to fine-tune a model’s parameters, denoted by , to accurately predict material compositions from a given sequence of elements. This process involves presenting the model with sequences of tokens , each representing either a unique element within a material composition or a numerical value representing the storage capacity. The model learns by adjusting to maximize the log-likelihood of correctly predicting the next token in the sequence, formalized as . Here, represents the conditional probability of observing the next token given the preceding sequence . The self-attention mechanism within the model facilitates this learning process by allowing each token to dynamically influence the prediction of the next token based on its context and relevance to the rest of the sequence.
@@
Here the critical hydrogen chemical potential, is defined as , which is lower than the chemical potential of the hydrogen at gas phase at very low temperature (as our calculation corresponds to DFT temperature scale). ”FE” stands for the formation energy. From the above table, it is apparent that our model exhibits commendable performance and is quite reliable for the calculation of the hydrogen storage capacity. In the final table 2, we present an additional five materials that have been predicted by GPT-2 and whose hydrogen weight percentage has been computed by our model. As the atomic configurations of these materials remain elusive, we employ the methodology expounded by Kusaba et al 58, which employs the approach of metric learning, where the algorithm has been trained on a substantial amount of crystal structures that have already been identified to ascertain the isomorphism of crystal structures formed by two specified chemical compositions, yielding an accuracy of approximately 96.4%. For a given composition with an unknown crystal structure, this approach automatically selects a collection of template crystals from a database of crystal structures that possess nearly identical stable structures. When we subject the compositions predicted by GPT-2 to this method, we obtain a collection of structures for each composition. We consider the one with the highest spacegroup and then perform the structural optimization followed by the calculation of the storage capacity using our hybrid-ML approach. The rightmost column of the table visually represents the crystal structure of each material.

### v4

@@ Abstract
In this study, we present a sophisticated hybrid machine-learning framework that significantly improves the accuracy of predicting hydrogen storage capacities in metal hydrides. This is a critical challenge due to the scarcity of experimental data and the complexity of high-dimensional feature spaces. Our approach employs the power of unsupervised learning through the use of a state-of-the-art autoencoder. This autoencoder is trained on elemental descriptors obtained from Mendeleev software, enabling the extraction of a meaningful and lower dimensional latent space from the input data. This latent representation serves as the basis for our deep multi-layer perceptron (MLP) model, which consists of five layers and shows good precision in predicting hydrogen storage capacities. Furthermore, our results show very good agreement with the results obtained with density functional theory (DFT). In addition to addressing the limitations caused by limited and unevenly distributed data in the field of hydrogen storage materials, we also focus on discovering new materials that show promising opportunities for hydrogen storage. These materials were identified using both feature-based approaches and predictions generated by a large language model. Finally, our investigation into the effectiveness of transferring weights from the autoencoder to the MLP, in addition to the latent features, suggests that while this strategy slightly improves model performance as indicated by a slightly higher R2 value and lower RMSE, it emphasizes the intricate challenge of adapting pre-trained weights for specific supervised tasks.
@@
The introduction of data-driven materials design 25, 26, 27 has accelerated material discovery, processing, and manufacturing 28, 29, 30. Machine learning (ML) algorithms allow for the creation of new alloys based solely on previously collected data, either publicly available or reported in scientific literature. Metal hydrides are an ideal subject for machine learning algorithm research due to their diverse properties and broad range. Traditional methods for determining the optimized materials class and corresponding metal hydride composition based on desirable properties are difficult, time-consuming, and costly. ML learning techniques, on the other hand, enable rapid, productive, and efficient material class prediction for a specific hydrogen weight percent and operational conditions. Numerous groups have begun to use ML techniques to speed up this screening and gain more physical insight from massive amounts of data 31, 32, 33. Rezakazemi et al. used an adaptive neuro-fuzzy inference system to evaluate the performance of hydrogen-selective mixed matrix membranes under various operational conditions 34. Rahnama et al. predicted the hydrogen storage capacities in various metal hydrides using machine learning algorithms in another study and found that higher temperatures yielded higher hydrogen storage capacities 33, 35. Ahmed et al. recently predicted gravimetric and volumetric hydrogen capacities in metal-organic frameworks (MOFs) using ML algorithms 36. ML methods have been instrumental in predicting various properties of (MOFs) and even designing new structures. This approach has significantly changed the research landscape in this field, allowing for the rapid identification of promising MOF candidates 37, 38 for various applications, including hydrogen storage. In a recent wor, Sun et al. introduced a method that combines meta-learning and high-throughput molecular simulations to predict hydrogen storage in nanoporous materials efficiently. 39
@@ 2 Methods
The architecture of our method is shown in Fig. 1, in a schematic way. We use a hybrid approach that has two parts: in the unsupervised learning part, an autoencoder is trained on the entire dataset, which consists of only unlabeled data, to learn a compact representation of the data in the bottleneck layer 40. A Multi-Layer Perceptron (MLP) is used in the supervised learning part to predict the hydrogen storage capacity. The major use of such an approach is to address the challenges of data scarcity and high feature dimensionality. In this study, the initial feature space has a high dimensionality and the available dataset is relatively small. As there are not too many machine learning studies available in the literature, it is not easy to choose proper descriptors. We have constructed 36 features using the Mendeleev software 41. To handle high-dimensional feature spaces with small datasets, we used an autoencoder to reduce the dimensionality of the data.
@@ 2.1 Unsupervised learning
Autoencoders are a kind of neural network that is used to extract features in an unsupervised manner. They are divided into two parts: an encoder, which compresses the input data into a lower-dimensional representation (also known as the latent representation or bottleneck), and a decoder, which reconstructs the original input from the compressed form 42.
@@
where is the activation function, and are the weight and bias parameters of the encoder, respectively. For the activation function, we have used the RelU function as implemented in PyTorch 43.
@@
This loss function was minimized using the Adam optimization method to find the optimal values for the parameters and say and 44. Once the autoencoder is trained, the encoder function can be used to extract features from the input data by using the activations of the lower-dimensional representation as the features.
The network architecture for both encoder and decoder includes two linear layers and dropout layers for preventing overfitting. The input to the autoencoder is first passed through a linear layer with 256 units, followed by a Leaky ReLU activation function 45. This is then followed by a dropout layer with a dropout probability of 0.3 to provide regularization.
@@ 2.2 Supervised learning
First-principles calculations were conducted employing the density functional theory (DFT) within the Vienna Ab initio Software Package (VASP) 46, 47. To determine the electronic energy, projected augmented wave (PAW) potentials and plane-wave basis sets were utilized for both core and valence electrons, employing the Perdew–Burke–Ernzerhof (PBE) functional 48, 49. The calculations accounted for long-range interactions using the semi-empirical Grimme D2 dispersion method and incorporated non-spherical contributions to the PAW potentials within the code 50. A cutoff energy of 500 eV ensured convergence of all energies. Structural optimization employed the conjugate gradient algorithm 51, with convergence criteria set at 10-7 eV for energy and 0.01 eVÅ-1 for force. Spin polarization was enabled in all calculations, except for the isolated H2 closed shell molecule.
@@ 3 Dataset
For this study, we utilized the Hydrogen Storage Materials Database shared by the HyMARC Data Hub (https://datahub.hymarc.org/). This is an openly available database that can be accessed at https://datahub.hymarc.org/en/dataset/hydrogen-storage-materials-db. This database contains information on more than 2000 hydride materials and their properties. The Hydride Database classifies the hydrides into eight classes: A2B intermetallic compounds, AB intermetallic compounds, AB2 intermetallic compounds, AB5 intermetallic compounds, complex hydrides, Mg alloys, solid solution alloys, and Misc. In the dataset, more than 950 compounds have 1-2 wt% hydrogen storage capacity (Fig. 2). The average hydrogen storage capacity of all compounds is 2.1 wt%. The dataset has as high as 20.8 wt% hydrogen storage capacity complex hydrides, which is BeB4H8 52.
@@ 5.2 Case I: Using the MLP Directly for All the Features
For our primary experiment, the MLP was trained and tested using all of the 36 features derived from the Mendeleev software. The dataset was strategically partitioned into training and testing subsets, with respective ratios of 0.8 and 0.2. The resultant outcomes of this methodology are visually represented in Fig.4. The predictive efficacy, quantified by the coefficient, was deemed unsatisfactory 53. Such results emphasize the inherent difficulties connected with managing datasets with a high number of dimensions, especially when the amount of data is small.
@@ 6 Comparison with few other supervised models
In this section, we compare the performance of two supervised models, Lasso regression and Gradient Boosting regression, with the previously mentioned hybrid approach using an Autoencoder combined with a Multilayer Perceptron (MLP) 54, 55. We evaluate the models based on their R2 scores, which measure the proportion of the variance in the target variable that the models can explain. The results are shown in Fig.7. The R2 values obtained from the Lasso regression and Gradient Boosting regression are 0.48 and 0.8, respectively. These values indicate that both models have relatively lower predictive accuracy compared to the hybrid approach. The hybrid approach achieved an R2 score of 0.85, outperforming both linear models. These results suggest that the hybrid approach, combining an Autoencoder with an MLP, is more effective in capturing the underlying patterns and predicting the target variable compared to the individual linear models.
@@ 6.1 Identification of new hydrogen storage materials
To find new materials for efficient hydrogen storage, we used the following approach: We calculated the Pearson correlation function for different physically meaningful features, including the target - the storage capacity (Wt), which we show in Fig. 8 56. We take the weighted sum for all nine features and compute the correlation. It can be seen that the period number and electronegativity have the highest correlation with the weight percentage. A significant Pearson correlation between the period number and electronegativity with the target hydrogen storage capacity, suggests their pivotal roles in influencing a material’s ability to store hydrogen efficiently. The period number, reflecting the electron shell structure, directly impacts its hydrogen bonding capabilities, while electronegativity, indicative of an electron attraction strength of the material, influences the stability and capacity of hydrogen storage in metal hydrides. Therefore, we first scanned the metal hydrides from the Materials Project with these two features. Next, we constructed a 36-dimensional feature space using the above materials, as before (using four statistical measures for each one). Then, we used pre-trained embeddings from the autoencoder so that the property (hydrogen weight in percentage) can be computed using the MLP. Finally, we performed DFT-based calculations to re-verify these values.
Using the above approach we identified two materials Al11O18 and V2O5. Both the materials are stable from our formation energy calculations. The DFT calculated formation energies are respectively -3.34 eV/atom and -2.12 eV/atom for Al11O18 and V2O5. From the DFT approach the hydrogen storage capacities were calculated using the hydrogen adsorption energies 11 shown below,
@@
To predict hydrogen storage capacity, all interstitial sites of Al11O18 and V2O5 are gradually filled with hydrogen, and the adoption energy is calculated 57. Figure 9 shows the change in adsorption energy with increasing hydrogen concentration. The adsorption energy increases as the hydrogen fraction increases. The maximum hydrogen fraction at which adsorption energy remains negative is used to predict hydrogen storage capacity. The predicted hydrogen storage capacities from the DFT calculations are respectively 4.61% and 3.83% while from the MLP model, we obtain values respectively 5.98% and 4.35%.
In the Fig.10, we show the density of states (DOS) for the machine learning predicted materials and their hydrides for different hydrogen loadings. It was found that Al11O18 shows a more stable behavior when loaded with hydrogen compared to V2O5 because V2O5 relatively DOS at the Fermi energy for all the hydrogen loading. For both materials, H-s strongly hybridizes with the O-p states as can be seen from the projected DOS of Al11O18H4 and V2O5H shown in the down panel of the above figure. In the case of V2O5H, the anti-bonding states formed between V-d and H-s are also filled and therefore the system shows less stability.
@@ 7 Application to the Hydrogen Storage Materials Generated Using a Large Language Model
To show further the predictive capability of our approach, we apply it further to a set of materials that we generate using a large language model (LLM). Leveraging the power of GPT-2 (Generative Pre-trained Transformer version 2), a model renowned for its proficiency in natural language processing, we have generated a few new hydrogen storage materials. Unlike the resource-intensive traditional methods such as Generative Adversarial Networks (GANs) or diffusion models, which have limited testing in materials science, our approach utilizes a fine-tuned GPT-2 model, offering a more quick and targeted pathway to innovation. The details of the methodology are explained as detailed by Fu et. al 58 which leverages the advanced capabilities of LLMs originally developed for natural language processing, adapting them to the intricate language of material compositions. The fine-tuning process involves adapting a pre-trained GPT-2 model to the specific nuances of material science. This adaptation is facilitated by a custom vocabulary. which includes a comprehensive list of chemical elements and special tokens. This specialized vocabulary is crucial for accurately representing and processing material compositions. Our training process fine-tunes the model with a dataset specifically curated for hydrogen storage materials, allowing the model to learn and predict the complex patterns of material composition. The cornerstone of this methodology is the training of these models on comprehensive databases of known material compositions, including the ICSD (Inorganic Crystal Structure Database), OQMD (Open Quantum Materials Database), and Materials Project. This extensive assimilation of existing material data equips LLMs like GPT-2 with the ability to understand and replicate the intricate patterns and rules that govern material compositions. Through this process, the models are enabled to generate new, chemically plausible material compositions, which hold great potential for hydrogen storage applications. In an extension to the work by Fu et al. 58 on materials generation, we take a slightly modified approach where not only the chemical composition of materials but also a specific property: their hydrogen storage capacity is also considered to generate the chemical labels.
@@
Starting from a sequence of input tokens (comprising chemical symbols and storage capacity) the model first computes their embeddings. It then applies learned weight matrices to derive the usual queries, keys, and values matrices , , and , which are used to calculate the self-attention scores given by 59 where is the dimensionality of the key vectors. These scores determine how much each token in the sequence should attend to every other token, capturing both local and global dependencies within the sequence. During the training phase, the objective is to fine-tune a model’s parameters, denoted by , to accurately predict material compositions from a given sequence of elements. This process involves presenting the model with sequences of tokens , each representing either a unique element within a material composition or a numerical value representing the storage capacity. The model learns by adjusting to maximize the log-likelihood of correctly predicting the next token in the sequence, formalized as . Here, represents the conditional probability of observing the next token given the preceding sequence . The self-attention mechanism within the model facilitates this learning process by allowing each token to dynamically influence the prediction of the next token based on its context and relevance to the rest of the sequence.
@@
Here the critical hydrogen chemical potential, is defined as , which is lower than the chemical potential of the hydrogen at gas phase at very low temperature (as our calculation corresponds to DFT temperature scale). ”FE” stands for the formation energy. From the above table, it is apparent that our model exhibits commendable performance and is quite reliable for the calculation of the hydrogen storage capacity. In the final table 2, we present an additional five materials that have been predicted by GPT-2 and whose hydrogen weight percentage has been computed by our model. As the atomic configurations of these materials remain elusive, we employ the methodology expounded by Kusaba et al 60, which employs the approach of metric learning, where the algorithm has been trained on a substantial amount of crystal structures that have already been identified to ascertain the isomorphism of crystal structures formed by two specified chemical compositions, yielding an accuracy of approximately 96.4%. For a given composition with an unknown crystal structure, this approach automatically selects a collection of template crystals from a database of crystal structures that possess nearly identical stable structures. When we subject the compositions predicted by GPT-2 to this method, we obtain a collection of structures for each composition. We consider the one with the highest spacegroup and then perform the structural optimization followed by the calculation of the storage capacity using our hybrid-ML approach. The rightmost column of the table visually represents the crystal structure of each material.